Mike can paint a room in 14 hours Bob could paint the same room in 10 hours. If they work together how long would it take them to complete job.
Answer by nerdybill(7384)

Mike can paint a room in 14 hours Bob could paint the same room in 10 hours. If they work together how long would it take them to complete job.
.
Let x = time (hours) it takes for both
then
x(1/14 + 1/10) = 1
multiplying both sides by 140 (common denominator):
x(10 + 14) = 140
x(24) = 140
x = 140/24
x = 5.833 hours
or
x = 5 hours and 50 minutes